4. The table shows the height of a plant as it grows.
a. Model the data with an equation.
b. Based on your model, predict the height of the plant at 12 months.

Time (months) Plant Height (cm)
3 9
5 15
7 21
9 27

Answer by checkley77(12844)

9/3=3
15/5=3
21/7=3
27/9=3
THUS 3 IS THE MONTHLY MULTIPLIER.
12*3=36 IS THE PREDICTED HEIGHT OF THE PLANT AFTER 12 MONTHS.